OpenStudy (anonymous):

Which variable expression represents the word phrase? six decreased by the sum of a number and two A. (m + 2) – 6 B. 6 – (m + 2) C. 6 – (m – 2) D. (m – 2) + 6

OpenStudy (anonymous):

B. by inspection.

OpenStudy (anonymous):

the answer is b

OpenStudy (anonymous):

Is your answer a question or a statement?

OpenStudy (anonymous):

it is a question

OpenStudy (anonymous):

am taking a k12 test

OpenStudy (anonymous):

bohotness what do you think

OpenStudy (bohotness):

what grade?

OpenStudy (anonymous):

8th

OpenStudy (anonymous):

a b or c

OpenStudy (bohotness):

i think b ?

OpenStudy (anonymous):

Sorry. Let m be the number. six decreased by the sum of a number and two 6 - ( + m + 2 ) ; -> B.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

i have some more qestions

OpenStudy (anonymous):

go ahead.

OpenStudy (anonymous):

Which equation represents the word sentence? The product of 5 and a number increased by 7 is 32. A. 5x = 32 + 7 B. 5x – 7 = 32 C. 5x + 7 = 32 D. 5x + 32 = 7

OpenStudy (anonymous):

Have to bail out for dinner. Will return in about an hour.

OpenStudy (anonymous):

okay same here